OpenStudy (123ab456c):

The county animal park is shaped like a triangle. Two perpendicular sides of the park measure 2.3 miles and 3.4 miles. What is the area of the animal park? A. 3.91 mi² B. 5.7 mi² C. 7.82 mi² D. 11.4 mi²

OpenStudy (123ab456c):

A

OpenStudy (welshfella):

the area of the park = 1/2 * 2.3 * 3.4 square miles

OpenStudy (123ab456c):

is the answer B

OpenStudy (welshfella):

No you were write first time 1/2 * 2.3 * 3.4 = 3.91 The area of a triangle = 1/2 * base * height

OpenStudy (123ab456c):

is the answer B

OpenStudy (welshfella):

the answer is as above

OpenStudy (123ab456c):

the answer is A

OpenStudy (welshfella):

yes

OpenStudy (123ab456c):

OK THANKS

OpenStudy (welshfella):

yw

OpenStudy (123ab456c):

what does that mean yw

OpenStudy (welshfella):

you are welcome

OpenStudy (123ab456c):

oh